OpenStudy (anonymous):

William is six years older than Robert. In two years, William will be twice as old as Robert. Find their ages.

OpenStudy (anonymous):

will=6*rob will+2=rob*2 solve everything in terms of will So second equation becomes will=rob*2-2 then just graph and find intersection graph

OpenStudy (anonymous):

I mean solve everything in terms of rob lol sorry

OpenStudy (anonymous):

also first line should be will=6+rob not 6*rob

OpenStudy (anonymous):

I'm confused....

OpenStudy (anonymous):

you can also plug in will=rob*2-2 into will=6+rob it becomes rob*2-2=6+rob

OpenStudy (anonymous):

rob should be 8

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

will=6+rob will+2=rob*2 since rob is 8 you can plug that in for rob will= 6 + 8 = 14 will is 14

OpenStudy (anonymous):

Okay I've got an equation. Let's say Williams age is W and Robert's is R... (6 + R) + 2 = 2(R +2) I think that'll help us both. xD

OpenStudy (anonymous):

William is six years older than Robert. w=6+R In two years, William will be twice as old as Robert. Find their ages. w*2=2+R

OpenStudy (anonymous):

Okay I've got an equation. Let's say Williams age is W and Robert's is R... (6 + R) + 2 = 2(R +2) I think that'll help us both. xD In your equation rob is 4 plug that back in as robert being 4 for w=6+R w=6+4 10 I think :)

OpenStudy (anonymous):

yes! :DDD I think we've got it!